OpenStudy (anonymous):

find k so that the line through (5,4) and (k,7) is a) parellel to 4y-3x=7 b) perpendicular to 2y+5x=1

OpenStudy (lgbasallote):

parallel = same slope...the slope is 4y = 3x+7 so slope = 3/4 k-4/7-5 = 3/4 k-4/2 = 3/4 cross multiply k-4 = 3/2 k = 3/2 + 8/2 k = 11/2